### v2

# Intelligent Graybox Fuzzing via ATPG-Guided Seed Generation and Submodule Analysis

Abstract
Hardware Fuzzing emerged as one of the crucial techniques for finding security flaws in modern hardware designs by testing a wide range of input scenarios. One of the main challenges is creating high-quality input seeds that maximize coverage and speed up verification. Coverage-Guided Fuzzing (CGF) methods help explore designs more effectively, but they struggle to focus on specific parts of the hardware. Existing Directed Gray-box Fuzzing (DGF) techniques like DirectFuzz try to solve this by generating targeted tests, but it has major drawbacks, such as supporting only limited hardware description languages, not scaling well to large circuits, and having issues with abstraction mismatches. To address these problems, we introduce a novel framework PROFUZZ that follows the DGF approach and combines fuzzing with Automatic Test Pattern Generation (ATPG) for more efficient fuzzing. By leveraging ATPG’s structural analysis capabilities, PROFUZZ can generate precise input seeds that target specific design regions more effectively while maintaining high fuzzing throughput. Our experiments show that PROFUZZ scales 30× better than DirectFuzz when handling multiple target sites, improves coverage by 11.66%, and runs 2.76× faster, highlighting its scalability and effectiveness for directed fuzzing in complex hardware systems.
I Introduction
Modern hardware systems, from resource-constrained compact edge devices to complex, heterogeneous platforms, are rapidly evolving to support the increasing demands of next-generation applications. These designs often integrate billions of transistors and a wide range of components, making the hardware verification process both critical and increasingly challenging. Despite significant advances, traditional verification techniques, including functional and formal methods [1, 2, 3], struggle to scale with the growing complexity of hardware designs to expose subtle or potential vulnerabilities. Recent advancements in leveraging Large Language Models in automating bug identification and mitigation [4, 5, 6, 7, 8, 9, 10] have shown promise, but these techniques remain limited in their scope and accuracy, paving the way for fuzzing as a complementary approach to uncover deeper, hard-to-find bugs. Fuzzing is an automated testing method that repeatedly generates inputs, often randomly or through guided mutation, to expose bugs, crashes, or unexpected behaviors [11]. Inspired by well-established software fuzzing techniques, hardware fuzzing techniques emerged as a promising verification approach [12, 13, 14, 15, 16, 17, 18, 19, 20]. Coverage-Guided Graybox Fuzzing (CGF) [12] has become a foundational technique in this space, and RTL-level Fuzzing frameworks like [18, 21, 14, 22] have demonstrated effectiveness in achieving broad coverage, particularly across designs with complex datapaths [23, 24]. In CGF, input generation is directed by feedback from real-time code coverage, allowing the fuzzer to prioritize input mutations that reveal unexplored paths in the design to improve functional coverage of RTL designs. While CGF-based techniques mark a significant advancement, they fall short in addressing the practical realities of hardware design workflows. Unlike software, hardware is typically developed incrementally, with new modules added or existing ones modified across design iterations. In such flows, verification efforts must focus on specific changes rather than re-testing the entire system. Directed Graybox Fuzzing (DGF) has been proposed as an alternative, which, unlike CGF, enables targeted fuzzing by steering test generation toward specific regions of interest. This is particularly useful in scenarios such as patch validation, bug localization, or module-specific testing. DirectFuzz [25], a DGF-based implementation, aimed to bring directionality into the fuzzing process to improve targeted coverage. Despite its innovation, DirectFuzz exhibits several key limitations, including (i) inability to capture hardware-specific semantics and structure, (ii) misaligned coverage metrics that do not align well with established verification practices, (iii) limited scalability when targeting multiple design regions, and (iv) an inability to perform targeted verification across multiple modules. These challenges motivate the need for a more scalable and structurally aware directed fuzzing framework tailored to the realities of modern hardware design and verification.
In this work, we present PROFUZZ, a novel hardware fuzzing framework based on DGF, specifically designed to overcome the challenges, such as abstraction mismatches, limited coverage precision, and poor scalability. Unlike existing methods, PROFUZZ operates directly at the hardware’s native abstraction level, allowing it to accurately model inherent hardware behaviors and utilize coverage metrics that reflect hardware-specific characteristics. Additionally, our framework seamlessly integrates with industry-standard Electronic Design Automation (EDA) tools, facilitating easy adoption within established design and verification workflows.
Our main contributions are summarized as follows:
  • Hardware-native Directed Graybox Fuzzing framework: Captures intrinsic hardware characteristics and delivers efficient coverage scaling with increasing complexity of target regions.
  • Scalable target site selection strategy: Supports both random and cost-driven approaches, adaptable to diverse hardware designs.
  • Integration of Automatic Test Pattern Generation (ATPG): Enables directed seed generation for target sites, utilizing conflict-aware merging to enhance seed quality and mutation diversity.
  • Cross-module verification support: Employs automated submodule generation to enable comprehensive fuzzing across multiple design regions.
II Background
II-A Fuzzing in Hardware: CGF and DGF
Hardware fuzzing [18, 15, 21] is a dynamic testing technique that exposes hidden bugs and vulnerabilities by subjecting the Design Under Test (DUT) to randomized input stimuli. Instead of relying purely on random inputs, most hardware fuzzers employ CGF, which uses coverage feedback to steer input mutations and efficiently explore the design’s state space, as illustrated in Fig. 1. The process starts with generating an initial set of input seeds, which are then iteratively mutated based on real-time coverage data collected from the DUT. Popular mutation strategies, borrowed from the American Fuzzy Lop (AFL) framework [26], include bit-flipping, byte swapping, and arithmetic operations. These techniques systematically modify inputs to exercise diverse execution paths and trigger rare hardware states. RFUZZ [12] is among the pioneering hardware fuzzing frameworks that apply CGF to maximize code coverage in hardware designs. By using coverage feedback, RFUZZ iteratively mutates inputs to accelerate coverage growth. However, a broad exploration strategy of CGF-based techniques often results in inefficient use of computational resources. More importantly, it lacks the targeted focus necessary to validate specific modules or recent incremental changes, limiting its applicability for hardware development workflows that rely heavily on iterative design updates.
Figure 1: Overview of Coverage Graybox Fuzzing (CGF).
In contrast, DGF focuses on specific target sites within hardware designs by monitoring coverage at these sites during fuzzing and applying mutations to input seeds based on improvements in targeted coverage, as illustrated in Fig. 2. The first practical implementation of DGF for hardware, DirectFuzz [25], aims to maximize coverage in designated regions of a design. It achieves this by translating RTL code into a software model using Verilator [27], treating multiplexer selection signals similarly to software branches to serve as coverage points. DirectFuzz further builds a module connectivity graph to steer input generation toward activating these signals. While this approach effectively targets specific hardware sites, it has notable drawbacks, especially its dependence on software-level abstractions and limited scalability.
Figure 2: Overview of Directed Graybox Fuzzing (DGF).
TABLE I: Comparative analysis of the proposed PROFUZZ approach with existing CGF and DGF solutions
Technique Abstraction Method Input Simulator
Iterative
Approach?
#Target Sites
(Min - Max)
EDA Toolflow
Integration?
[…]
DirectFuzz [25] RTL (SW) DGF Series of bits Verilator 6 - 107
[…]
$\blacklozenge$
any: Any commercial industry-standard HDL simulator.
II-B Categorizing Hardware Fuzzing Frameworks
Existing hardware fuzzing frameworks can be broadly classified into three categories:
Direct Use of Software Fuzzers on Hardware
This method applies software fuzzers like AFL [26] directly to hardware designs, generating randomized inputs and using software-oriented coverage metrics (e.g., branch coverage) to guide the fuzzing process.
Fuzzing Hardware as Software
Hardware designs written in RTL languages such as Verilog or VHDL are translated into software models using tools like Verilator. These software abstractions are then fuzzed with standard software fuzzers.
Fuzzing Hardware as Hardware
This approach performs fuzzing directly on hardware representations, such as RTL or gate-level netlists, leveraging industry-standard EDA tools like Synopsys VCS [28] or Cadence Xcelium [29]. These tools provide native hardware coverage metrics, including branch, toggle, FSM, expression, and line coverage, aligned with conventional hardware verification flows.
II-C Prevailing Challenges of Directed Gray-box Fuzzing
II-C1 Software-Level Abstraction
Translating RTL to C++ through Verilator creates a mismatch by failing to capture key hardware semantics like concurrent execution, precise timing constraints, register behavior, and signal interactions. This abstraction gap means the software model doesn’t fully represent the actual hardware behavior, limiting the effectiveness of fuzzing based on it.
Challenge C1: The translated software version of the hardware by Verilator fails to account for the inherent hardware characteristics.
II-C2 Misaligned Coverage Metrics
DGF relies on coverage feedback derived from the software model generated by Verilator rather than from the hardware itself. While some software constructs like branches loosely correspond to hardware elements such as multiplexers, this mapping is limited. Important hardware coverage aspects, like signal toggling, FSM transitions, expression evaluation, and block activation, are either lost or poorly represented. This abstraction gap leads to coverage metrics that do not accurately reflect true hardware behavior, causing the fuzzing process to focus on inputs that improve software-like coverage while missing other critical hardware states.
Challenge C2: The coverage metric fails to capture intrinsic hardware characteristics associated with the actual hardware design.
II-C3 Lack of Scalability
Current DF approaches are designed to handle <<100 target signals limited to specific structures like multiplexers (MUX), which limits the scalability as target regions grow. It also lacks flexibility to handle important control or datapath signals beyond MUX logic, restricting its ability to support broader and more complex verification tasks.
Challenge C3: Current DGF techniques like DirectFuzz target fewer than 100 signals, limiting their scalability and failing to cover broader target regions.
II-C4 Lack of Cross-Module Verification
Current DFG approaches like DirectFuzz [25] focus on individual modules, such as targeting only the RX or TX block in a UART design. While effective for isolated testing, this approach falls short when verifying interactions across multiple modules. Since many real-world bugs stem from inter-module dependencies, the inability to validate cross-module behavior is a critical limitation for comprehensive hardware verification.
Challenge C4: Current DGF techniques like DirectFuzz restrict the fuzzing to single target module at a time, making it ineffective for verifying interactions across multiple interconnected components.
Table I summarizes existing hardware fuzzing approaches, categorizing them by abstraction level and core methodology, while highlighting the need for our proposed approach to address the limitations of current techniques.
Figure 3: Overview of PROFUZZ methodology highlighting major steps.
III PROFuzz Methodology
Fig. 3 demonstrates the overview of the PROFUZZ methodology, highlighting three major components: (1) Target Site Identifier, (2) Seed Generator, and (3) Directed Fuzzing. The input to the Target Site Identifier is the hardware design ($\mathcal{I_{N}}$) at its native abstraction level (RTL or gate-level). PROFUZZ constructs a hypergraph $G=(\mathcal{V},\mathcal{E})$ from the netlist abstraction (for RTL, it performs initial synthesis to generate the gate-level netlist), where $\mathcal{V}$ represents the set of gates and $\mathcal{E}$ the interconnections between them. Using configuration parameters ($\theta$), the Target Site Identifier selects a subset of nodes ($\mathcal{T_{N}}$) for verification and coverage improvement. Next, PROFUZZ applies conflict-PI (Primary Input) and fan-in cone analysis to identify non-conflicting nets and leverages ATPG to generate activation patterns. These patterns are merged to create a targeted seed pattern ($\mathcal{P_{T}}$) that maximizes mutation potential. The directed fuzzing loop then iteratively mutates inputs, monitors coverage at target sites ($\mathcal{C_{T}}$), and updates the seed pool until a coverage goal or timeout is reached. The following sections provide a detailed overview of each stage.
III-A Target Site Identifier
PROFUZZ generates seed inputs to maximize coverage of target sites in hardware designs (Fig. 3). Unlike DirectFuzz, which uses signals from C++-translated models, PROFUZZ operates directly on native hardware abstractions, preserving design semantics and addressing challenges C1 and C2. Target signals ($\mathcal{T_{N}}$) can be selected manually or automatically. In the manual approach, the verification engineer selects signals based on specific verification requirements. For automatic selection, a cost function based on structural (fan-in $\mathbb{FI}$, fan-out $\mathbb{FO}$) and stochastic properties (Shannon entropy $\mathbb{H}$) is used to select target signals. This function is applied to each edge $i\in\mathcal{E}$ as follows:
$\mathcal{C}_{i}={cost\_fun(\mathbb{FI}_{i},\mathbb{FO}_{i},\mathbb{H}_{i})},\forall i\in\mathcal{E}$ (1)
The nets/signals are ranked and selected based on $\mathcal{C}_{i}$ values along with heuristics such as controllability, observability, and topological depth to maximize verification efficiency.
Input : Gate-level Netlist ($\mathcal{I_{N}}$), config ($\theta$)
Output : Simulation Vectors ($\mathcal{S_{I}}$), Target Site Coverage ($\mathcal{C_{T}}$)
1
2GG $\leftarrow$ parse $\mathcal{I_{N}}$ and construct hypergraph $G=\{\mathcal{V},\mathcal{E}\}$
3 $G^{T}$ $\leftarrow$ topological_sort($G$)
/*        Target Site Identifier     */
4 $\mathcal{T_{N}}$ $\leftarrow$ target_net_selection($G,\theta$)
/*         Seed Generator     */
5 $\mathcal{P_{T}},\mathcal{N}_{0},\mathcal{N}_{1}$ $\leftarrow$ pattern_generation($G,G^{T},\mathcal{T_{N}},\theta$)         //       $\triangle$ Detailed in Algo 2
6 $\mathcal{S_{T}}$ $\leftarrow$ $\mathcal{P_{T}}$, $\mathcal{C_{T}}\leftarrow 0$, $\mathcal{C_{T}^{\text{init}}}\leftarrow 0$
/*         Directed Fuzzing     */
7
8while termination criteria not met do
9 foreach seed $S\in\mathcal{S_{T}}$ do
10
11    $\mathcal{S_{M}}\leftarrow$ mutatemutate($S$)
12    $\mathcal{C_{T}}\leftarrow$ runrun($\mathcal{I_{N}}$, $\mathcal{S_{M}}$)
13
14    if $\mathcal{C_{T}}>\mathcal{C_{T}^{\text{init}}}$ then
15       $\mathcal{S_{I}}\leftarrow\mathcal{S_{I}}\cup\{S_{M}\}$
16       $\mathcal{S_{D}}\leftarrow\mathcal{S_{M}}$
17       $\mathcal{C_{T}^{\text{init}}}\leftarrow\mathcal{C_{T}}$
18
19
20 until time or coverage threshold
21return $\mathcal{S_{I}},\mathcal{C_{T}}$
Algorithm 1 PROFUZZ Methodology
Existing approaches, as noted in Section II-C, struggle with scalability when handling many target signals, while selecting too many can add significant overhead, resulting in degraded performance. To address this, PROFUZZ creates a lightweight target submodule that includes only the selected signals. This submodule is integrated into the original design, enabling precise monitoring without full DUT instrumentation, thus solving Challenge C3. This method reduces overhead and supports cross-module verification, addressing Challenge C4. The parameter ($\theta$) in Algorithms 1 and 2 configures the selection. Signals within the target submodule are labeled ‘target’ and the ATPG engine focuses on maximizing coverage for these nodes ($\mathcal{T_{N}}$).
Input : $G,G^{T},\mathcal{T_{N}},\theta$
Output : $\mathcal{P_{T}},\mathcal{N}_{0},\mathcal{N}_{1}$
1 $ID_{N}$ $\leftarrow$ get_id($G,G^{T},\mathcal{T_{N}}$)
2 $\mathcal{P}_{all}$ $\leftarrow$ gen_non_conflict_patt($G,\mathcal{T_{N}},ID_{N}$)
3 $\mathcal{P}_{merged},\mathcal{N}_{0},\mathcal{N}_{1},\mathcal{N}_{C}$ $\leftarrow$ merge_pattern($\mathcal{P}_{all}$)
4 $\mathcal{T_{N}}\leftarrow\mathcal{T_{N}}-\mathcal{N}_{C}$
5 if $\theta.submodule$ is TrueTrue then
[…]
8 $\mathcal{P}_{merged}^{M},\mathcal{N}_{0},\mathcal{N}_{1},\mathcal{N}_{C}$ $\leftarrow$ merge_pattern($\mathcal{P}_{all}^{M}$)
[…]
10 $\mathcal{P_{T}}\leftarrow\mathcal{P}_{merged}^{M}$
11 return $\mathcal{P_{T}},\mathcal{N}_{0},\mathcal{N}_{1}$
[…]
13 return $\mathcal{P_{T}},\mathcal{N}_{0},\mathcal{N}_{1}$
Algorithm 2 Pattern Generation
III-B Seed Generator
After identifying the target nets ($\mathcal{T_{N}}$), PROFUZZ generates seed patterns to maximize coverage of target sites during fuzzing. It integrates Synopsys TestMAX ATPG to create hierarchical IDs and activation patterns (0 and 1) for each net $n_{i}\in\mathcal{T_{N}}$ using the stuck-at-fault model. PROFUZZ performs conflict-PI and fan-in-cone analysis to identify conflicting nets ($\mathcal{N}C$) and merges non-conflicting patterns into a single seed pattern ($\mathcal{P}{merged}$). The merging maximizes Don’t Care (X) values for Primary Inputs (PIs), increasing the number of potential mutations. Based on the merged patterns, nets are assigned to $\mathcal{N}_{0}$ (for 0) or $\mathcal{N}_{1}$ (for 1). For target submodules, PROFUZZ iteratively generates seed patterns for both the submodule and the main module. Algorithm 2 outlines this ATPG-based pattern generation.
III-C Directed Fuzzing
The input design ($\mathcal{I_{N}}$) is fuzzed using targeted seed patterns ($\mathcal{P_{T}}$), forming the initial seed pool ($\mathcal{S_{T}}$) to stimulate the target submodule. As shown in Fig. 3, the coverage points correspond to the target nodes ($\mathcal{T_{N}}$), which represent the signals selected for focused verification. To track fuzzing effectiveness, PROFUZZ uses industry-standard EDA tools to analyze the design and measure coverage ($\mathcal{C_{T}}$) at these target nodes. The fuzzing process follows an iterative feedback loop: mutations are applied to each seed to generate new test vectors ($\mathcal{S_{M}}$), and only mutations that increase coverage are retained to update the seed pool. We use the AFL-style mutation algorithm [12]. The fuzzing campaign ends when a predefined coverage threshold is met or the allocated time expires, ensuring efficiency and controlled resource usage.
Figure 4: Integrating PROFUZZ into the commercial EDA tool flow.
III-D Integration into commercial EDA tool flow
PROFUZZ is fully compatible with the standard commercial EDA tool flow for ASICs and can be seamlessly integrated into existing verification and simulation environments for performing directed fuzzing for any hardware design. This compatibility ensures that PROFUZZ can be readily applied to a wide range of hardware designs to enhance functional coverage and uncover potential implementation bugs. Fig. 4 demonstrates the integration of PROFUZZ into the commercial (Synopsys/Cadence) EDA tool flow.
IV Experimental Evaluation
IV-A Experimental Setup
PROFUZZ is fully compatible with the industry-standard IC design and verification flow, enabling seamless Directed Fuzzing. For hardware design simulations, we use Cadence Xcelium, a widely adopted EDA tool in both industry and academia. Target state and site coverage metrics are extracted using Cadence Integrated Metrics Center (IMC) through custom TCL and Python scripts. Seed patterns for activating target sites are generated using Synopsys TestMAX (V-2023.12-SP5). Experiments were conducted on an Intel Core i7 (2.4 GHz) running a Linux OS. The fuzzing campaign continued until coverage reached $\geq$ 90% (configurable threshold).
IV-B Evaluation Dataset
We evaluate PROFUZZ across a diverse set of hardware designs as illustrated in Table II. For each of the designs, we assess the effectiveness of our proposed framework by performing target site selection under two configurations: (1) random selection and (2) Cost Function-based selection. For the cost function-based configuration, we selected target nodes using two criteria: (i) nodes with a cost function value $\mathcal{C}_{i}$ greater than 0.3 and (ii) nodes within the top 10 percentile of the ranked distribution.
We evaluate PROFUZZ across a range of hardware designs, as shown in Table II. For each design, we assess its effectiveness by performing target site selection in two configurations: (1) random selection and (2) Cost Function-based selection. In the cost function-based configuration, target nodes are selected based on two criteria: (i) nodes with a cost function value $\mathcal{C}_{i}$ >> 0.3, and (ii) nodes within the top 10 percentile of the ranked distribution.
TABLE II: Details of evaluation benchmarks
[…]
ORCTRL 108 3328 3383 Control Unit of OR1200 CPU
GPIO 70 905 834 General Purpose Input/Output
FIR 33 3222 2521 Finite Impulse Response Filter (DSP)
IIR 33 2801 2545 Infinite Impulse Response Filter (DSP)
PWM 54 3454 3402 Pulse Width Modulation
SPI 102 7617 7447 Serial Peripheral Interface
I2C 18 2125 2121 Inter-Integrated Circuit
UART 28 1522 1369 Universal Async. Receiver/Transmitter
PicoRV32 101 53306 53215 RISC-V Processor
DFT 63 189567 187168 Discrete Fourier Transform (DSP)
IDFT 63 190361 187986 Inverse DFT (DSP)
DES3 235 3807 3630 Triple DES Encryption
SHA256 515 8045 7884 Secure Hash Algorithm 256-bit
[…]
$\blacklozenge$#Gates include the count of combinational and sequential logic gates. All the designs are synthesized using gscl45nm standard cell library.
IV-C Evaluation Results
Table III presents the coverage (in %) and the runtime (in seconds) for the evaluation benchmarks with randomly selected target signals. PROFUZZ demonstrates high efficiency on datapath-centric designs such as FIR and IIR, achieving 100% coverage while targeting only 32.5% and 15.1% of internal signals, respectively, in under 5 seconds. GPIO attains 99.39% coverage in 2.36 seconds despite a moderate target ratio (35.9%). For more complex control-heavy designs like ORCTRL, PROFUZZ achieves 94.54% coverage while targeting just 12.5% of internal nodes, demonstrating strong robustness. In contrast, designs such as PWM and SPI require over 50% of internal nodes to be targeted, achieving 53.24% and 75.63% coverage, respectively, reflecting the increased difficulty of covering dense target sets; nonetheless, SPI reaches 75.63% coverage in only 5.18 seconds. The UART design targets 64.6% of internal signals and achieves 93.81% coverage in 15.21 seconds, illustrating scalability to medium-complexity modules. Cryptographic cores (DES3, SHA256) achieve over 98% coverage within short runtimes, indicating strong observability despite algorithmic complexity. For I2C, with 1209 target nodes, the proposed framework achieves 94.10% coverage in approximately 240 seconds, further highlighting scalability of PROFUZZ. However, for large designs such as PicoRV32, DFT, and IDFT, full instrumentation incurs significant overhead ($\blacktriangle$) and resource usage (in Table III), limiting practicality. To address this, we employ a target submodule generation strategy, with results in Table IV, enabling scalable analysis of large-scale designs.
TABLE III: Evaluation for Random Selection of Target Signals
[…]
ORCTRL 415 94.54% 20.41
GPIO 325 99.39% 2.36
FIR 1049 100% 4.81
PWM 2008 53.24% 10.38
SPI 2857 75.63% 5.18
I2C 1209 94.10% 240.37
UART 983 93.81% 15.21
PicoRV32 5287 $\blacktriangle Ins Overhead\blacktriangle$
IIR 423 100% 4.31
DFT 9145 $\blacktriangle Ins Overhead\blacktriangle$
IDFT 9518 $\blacktriangle Ins Overhead\blacktriangle$
DES3 470 98.62 3.64
SHA256 993 98.09 10.09
[…]
$\nabla$
#Target Signals denotes the number of signals selected randomly from the design, ranging between 2%-50% of the total signals.
Table IV presents the coverage values (in %) for different benchmarks with target submodule configuration and respective time (in seconds), demonstrating the ability of PROFUZZ to isolate and verify signal groups that span multiple modules and instances. The instrumentation overhead for large designs, such as PicoRV32, is effectively mitigated using the target submodule generation strategy, which enables directed seed generation focused on specific regions, as presented in Table IV. For PicoRV32, with 455 target nodes, PROFUZZ achieves 60.42% coverage, while for ORCTRL it achieves similar coverage (with 415 nodes) under 20 seconds, but a more deeply embedded subset of 200 nodes requires nearly 300 seconds, highlighting the increased effort to activate nested logic. Smaller modules like GPIO, FIR, and I2C show consistent performance due to easier access to target sites. UART achieves 95.63% coverage for 219 target nodes, comparable to the full-module result of 93.81% with 983 targets (Table III), further validating the submodule approach.
TABLE IV: Evaluation for Generated Target Submodule
Design
#Sub_Target
[…]
#Main_Target
[…]
Coverage Time (s)
ORCTRL 200 749 92.23% 300.51
GPIO 416 209 99.94% 2.47
FIR 587 882 100% 2.63
PWM 710 975 55.38% 11.43
I2C 437 597 93.15% 240.37
UART 219 189 95.63% 2.59
DFT 307 12454 97.84% 60.71
IDFT 387 12749 96.86% 75.12
PicoRV32 455 5189 60.42% 636.24
MIPS 268 3129 52.48% 620.12
Sodor1Stage 707 2891 59.36% 660.57
Sodor3Stage 680 3003 61.27% 710.68
$\blacklozenge$#Sub_Target Signals: target signals encapsulated inside the target submodule.
$\blacktriangle$#Main_Target Signals: signals in the original design connecting the target submodule.
Fig. 5 shows the coverage (%) and number of target signals for various benchmarks selected using two different cost functions: (1) nets with $\mathcal{C}_{i}\geq 0.3$ and (2) nets in the top 10 percentile ($P_{10}$) of the ranked cost function distribution. PROFUZZ achieves over 90% coverage across all benchmarks except SPI. The number of target signals is also highlighted to demonstrate PROFUZZ’s scalability. Notably, coverage remains consistent across both cost function configurations, demonstrating the framework’s ability to effectively cover target nets ranked by their cost function values.
TABLE V: Comparing the performance of PROFUZZ with RFUZZ and DirectFuzz on various designs and highlighting the coverage improvement and speedup achieved by PROFUZZ over DirectFuzz
Benchmark (Target Inst.) #Signals Target Inst. Cell % RFUZZ[12] DirectFuzz[25] PROFUZZ (This work)
[…]
UART (Tx) 6 5.1% 100% 7.35 100% 0.42 100% 1.31 iso* 0.32
UART (Rx) 9 6.9% 88.89% 4.95 88.89% 1.71 100% 1.43 12.50% 1.20
SPI 5 34.4% 100% 55.84 100% 31.35 100% 3.51 iso* 8.93
PWM 14 26.9% 100% 12.79 100% 5.87 100% 2.76 iso* 2.13
FFT 107 87% 13% 0.075 13% 0.073 100% 3.47 669.23% 0.02
I2C 65 31% 98% 13.73 98% 8.49 100% 4.57 2.04% 1.86
Sodar1Stage (CSR) 93 16.6% 96.77% 500.56 96.77% 463.63 100% 210.51 3.34% 2.20
Sodar1Stage (CtlPath) 68 0.3% 100% 694.42 100% 526.53 100% 167.48 iso* 3.14
Sodar3Stage (CSR) 90 16.4% 98.89% 568.05 98.89% 446.29 100% 281.63 1.12% 1.58
Sodar3Stage (CtlPath) 66 0.3% 100% 1283.40 100% 1034.86 100% 236.94 iso* 4.37
Average 89.56% 314.12 89.56% 251.92 100% 91.36 11.66% 2.76
[…]
$\bigstar$
Speedup: Refers to the reduction in execution time, calculated as the ratio of the execution time of DirectFuzz to that of PROFUZZ.
Figure 5: Coverage(%) and the number of target signals (#Target Signals) for seven benchmarks with varying sizes where the target sites are selected based on cost functions: (1) $\mathcal{C}_{i}\geq 0.3$ and (2) $\mathcal{C}_{i}\in P_{10}$.
Figure 6: Whisker plots comparing the number of target sites handled by DirectFuzz and PROFUZZ.
IV-D Comparison with RFUZZ and DirectFuzz
Table V compares the performance of PROFUZZ with existing state-of-the-art techniques, including RFUZZ[12] and DirectFuzz[25]. As shown in previous subsections, PROFUZZ demonstrates superior efficacy by covering a significantly larger set of target signals compared to RFUZZ and DirectFuzz. While covering fewer target sites in smaller designs (e.g., UART) is easier, PROFUZZ performs better on more complex designs (e.g., Sodor5Stage). For a fair comparison, we used the same number of target signals as reported in DirectFuzz for each benchmark. Notably, RFUZZ and DirectFuzz struggle with FFT, where the number of target sites (107) is the highest among all tested designs. As seen in Table V, PROFUZZ achieves higher average coverage, with 11.66% improvement and a 2.76× speed-up over RFUZZ and DirectFuzz across designs of varying complexities.
Figure 7: Coverage convergence of PROFUZZ with varying number of target sites selected based on various configurations (random, cost function, targeted submodule) for benchmarks: (a) ORCTRL and (b) FIR Filter.
Fig. 6 compares the number of target sites handled by PROFUZZ and DirectFuzz across multiple benchmarks, showing a 2× to 30× increase in the number of target sites covered by PROFUZZ. This highlights the scalability and effectiveness of our approach, demonstrating its ability to handle a much larger number of target sites, particularly for complex hardware designs. As illustrated in Fig. 7, PROFUZZ demonstrates a steady improvement in target site coverage over time for all the configurations for ORCTRL and FIR Filter, highlighting the effectiveness of its directed fuzzing and efficient seed exploration in covering target sites.
V Conclusion
This work introduces PROFUZZ, a Directed Graybox Fuzzing framework that enhances scalability by targeting specific regions within a hardware design. Unlike DirectFuzz, which struggles with large sets of target sites, PROFUZZ supports signal-level activation across multiple modules and hierarchical regions, achieving 2×-30× more target sites covered with a 2.76× speedup over DirectFuzz. However, overall coverage depends on the nature of selected target sites, as some nets may hold constant logic values and cannot be toggled. Future research should focus on developing more intelligent, structure-aware mutation strategies specifically designed for hardware fuzzing, particularly for non-ISA-based inputs represented as raw bitstreams. Tailoring mutation engines to the hardware domain will be essential for enabling effective fuzzing across a broader range of complex hardware designs beyond traditional CPU cores or popular open-source projects.